# An Alignment Between the CRA’s Essential Requirements and the ATT&CK®’s Mitigations

### v1

@@ Abstract
Index Terms:
Security requirements, legal requirements, regulations, security countermeasures, evaluation, mapping, gaps
@@ III Results
The main conclusion to draw from the two figures is that the alignment is generally very good. Regarding the $1212$ collated non-functional requirements of the CRA and the $4444$ mitigations for enterprises, only three mitigations could not be mapped. These are the ATT&CK®’s mitigations to have a threat intelligence platform, to build an out-of-band communications channel for secure exchanges during incident management, as also recommended by other frameworks [30], and to provide training for users on matters such as phishing and social engineering in general. All three gaps can be seen to be on the organizational side. Also the CRA’s essential requirements about security testing and vulnerability coordination have been seen as being on an organizational side instead of the technical product-side [8]. Thus, in addition to the disagreements already noted, some ambivalence is present regarding functional and non-functional but technical requirements and organizational cyber security requirements.
@@ IV Conclusion
Regarding future research, (a) it seems sensible to further continue the alignment studies of cyber security standards. Such studies are important because following the upcoming harmonized standards will in most cases provide a presumption of conformity and compliance with the CRA, as is typical also in many other domains in the European Union [34, 35]. However, alignment evaluations are not sufficient alone; (b) once the standards have been made, also they should be evaluated, including with respect to their practical usefulness, preciseness and clarity [12, 36], rigor, scope, complexity, and other related evaluation criteria for standards. In a similar vein, (c) it seems that empirical evaluations are mostly absent with respect to the ATT&CK®’s usefulness for practitioners, including with respect to the mitigations offered. As was noted, the framework operates at a rather high abstraction level, which may—or may not—decrease their practical usefulness.

### v2

@@ III Results
The main conclusion to draw from the two figures is that the alignment is generally very good. Regarding the $12$ collated non-functional requirements of the CRA and the $44$ mitigations for enterprises, only three mitigations could not be mapped. These are the ATT&CK®’s mitigations to have a threat intelligence platform, to build an out-of-band communications channel for secure exchanges during incident management, as also recommended by other frameworks [30], and to provide training for users on matters such as phishing and social engineering in general. All three gaps can be seen to be on the organizational side. Also the CRA’s essential requirements about security testing and vulnerability coordination have been seen as being on an organizational side instead of the technical product-side [8]. Thus, in addition to the disagreements already noted, some ambivalence is present regarding functional and non-functional but technical requirements and organizational cyber security requirements.
@@ IV Conclusion
Regarding future research, (a) it seems sensible to further continue alignment studies with alternative cyber security frameworks [34] and particularly standards. Such studies are important because following the upcoming harmonized standards will in most cases provide a presumption of conformity and compliance with the CRA, as is typical also in many other domains in the European Union [35, 36]. However, alignment evaluations are not sufficient alone; (b) once the standards have been made, also they should be evaluated, including with respect to their practical usefulness, preciseness and clarity [12], rigor, scope, complexity, and other related evaluation criteria for standards. In a similar vein, (c) it seems that empirical evaluations are mostly absent with respect to the ATT&CK®’s usefulness for practitioners. As was noted, the framework operates at a rather high abstraction level, which may—or may not—decrease its practical usefulness.